For the following reaction, the initial concentrations of CO is 0.500M and Cl2 is 1.00 M. What are the equilibrium concentrations of each species at 400K? (show all work)

COCl2(g) ⇌ CO(g) + Cl2 (g) Kc = 0.680 at 400 K
Answer #1

cole, g) = CO(g) + 46,692 K2=0.680 Reverse the reaction CO(g) + CO₂ (g) È Coll (9) 0-680 1-47 ICE table for the reaction colg.X; = 1.92 ; 12 = 0.26 I cant be greater than loo because in that case (4 = 100-x ull ise negative that is not possible So,use comment section for the doubts and clarifications